OpenStudy (anonymous):

pls integral of xsin^22xdx

OpenStudy (apoorvk):

You mean: \[\int x.sin^2(2x).dx\]^This?

OpenStudy (anonymous):

yes i mean this

sam (.sam.):

\[\int\limits x \sin ^2(2 x) \, dx\] \[\int\limits x(sin(2x))^2 \, dx\] Use \(sin(2x)=2sin(x)cos(x)\), \[\int\limits x(2sinxcosx)^2 \, dx\] \[\frac{1}{2}\int\limits x (1-\cos (4 x)) \, dx\] Expand \[\frac{1}{2}\int\limits (x-x \cos (4 x)) \, dx\] Split \[\frac{1}{2}\int\limits x \, dx-\frac{1}{2}\int\limits x \cos (4 x) \, dx\] Integration by parts for xcos(4x) u=x dv=cos(4x) du=dx v=(1/4)sin(4x) \[-\frac{1}{8} x \sin (4 x)+\frac{1}{2}\int\limits x \, dx+\frac{1}{8}\int\limits \sin (4 x) \, dx\]

sam (.sam.):

Then for sin(4x) Let t=4x dt=4dx \[\frac{1}{32}\int\limits \sin (t) \, dt-\frac{1}{8} x \sin (4 x)+\frac{1}{2}\int\limits x \, dx\] Integrate

OpenStudy (anonymous):

how did you come up with 1/2 \[\int\limits_{}^{}x(1-\cos4x)dx\]

sam (.sam.):

\[\int\limits\limits x(2sinxcosx)^2 \, dx\] To \[\frac{1}{2}\int\limits\limits x (1-\cos (4 x)) \, dx\]?

OpenStudy (anonymous):

yes

sam (.sam.):

That's actually a trigonometry identity, \[\sin ^2(2 x)=\frac{1}{2} (1-\cos (4 x))\]

OpenStudy (anonymous):

ok